sphenix-tpc-l AT lists.bnl.gov
Subject: Sphenix-tpc-l mailing list
List archive
Re: [[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.17 at 9:30 AM ET
- From: "Hughes, Charles [PHYSA]" <chughes2 AT IASTATE.EDU>
- To: "Hughes, Charles [PHYSA]" <chughes2 AT IASTATE.EDU>, Thomas K Hemmick <Thomas.Hemmick AT stonybrook.edu>, David Baranyai <david.baranyai AT cern.ch>, Martin Purschke <purschke AT bnl.gov>, Joel Vasquez <jvasquez AT bnl.gov>, Stephen Boose <boose AT bnl.gov>, Seth Howell <seth.howell AT stonybrook.edu>, "sphenix-tpc-l AT lists.bnl.gov" <sphenix-tpc-l AT lists.bnl.gov>, Jin Huang <jhuang AT bnl.gov>, "divaldo95 AT gmail.com" <divaldo95 AT gmail.com>, Ross Corliss <ross.corliss AT stonybrook.edu>, Michael Lenz <mlenz AT bnl.gov>, Benjamin Kimelman <benjamin.kimelman AT vanderbilt.edu>, Luke Legnosky <luke.legnosky AT stonybrook.edu>, John Haggerty <haggerty AT bnl.gov>, Kevin Mandracchia <kmandracchia AT bnl.gov>, Yeonju Go <ygo AT bnl.gov>, Bob Azmoun <azmoun AT bnl.gov>, Damon Miraglia <dmiraglia AT bnl.gov>, Gabor David <gabor.david AT stonybrook.edu>, Takao Sakaguchi <takao AT bnl.gov>, "osbornjd91 AT gmail.com" <osbornjd91 AT gmail.com>, "Robert P. Pisani" <pisani AT bnl.gov>, Thomas K Hemmick <tkhemmick AT gmail.com>, Frank Toldo <fatoldo AT bnl.gov>, John Kuczewski <jkuczewski AT bnl.gov>, shana Prifte <sprifte AT bnl.gov>, "josborn1 AT bnl.gov" <josborn1 AT bnl.gov>, Kristina Finnelli <kfinnelli AT bnl.gov>, "balazs.ujvari AT cern.ch" <balazs.ujvari AT cern.ch>
- Subject: Re: [[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.17 at 9:30 AM ET
- Date: Sun, 16 Mar 2025 16:49:45 +0000
Once the gains have been declared balanced by @Thomas K Hemmick, we should resurrect the line laser and include it in the nightly data taking.
I meant to write resurrect the diffuse laser here.
Sent from my iPhone
On Mar 16, 2025, at 8:00 AM, Hughes, Charles [PHYSA] <chughes2 AT iastate.edu> wrote:
Hello All,
Let's meet MONDAY 03.17 at 9:30 AM ET (3:30 PM CET) to discuss ongoing operations/analysis/planning for the TPC. In particular:
HV: TPC has continued to make progress, running nightly and through the weekend with shifters recovering trips. Last week, Yeonju and @Luke Legnosky both contributed their gain data to @Thomas K Hemmick and he applied @Go, Yeonju's calculations to make the 1st (proper) iteration of gain balancing with our new HV system. On Friday/weekend the shifters took the gain scan data, which should help determine the optimal operation point (at least in the absence of B field). We can discuss this as well as Tom's weather adjustments.
LV/FEEs: This week, we may run some or all of the FEEs with an upgraded firmware version depending on @Kuczewski, John's schedule. Last week, we determined that the noise increase in NS 3 R1 cannot be remediated simply and so we will live with it. Finally, we should continue any discussion about wandering pedestals and producing a proper map of them. @Sakaguchi, Takaomade a change last week which eliminated some of them.
DAQ: @Huang, Jin was able to produce a working fix to our dirty memory issues from last week which has been deployed to all EBDC machines for 4-5 days now. We can discuss.
Line Laser: @Seth Howell, @Azmoun, Bob, and @Thomas K Hemmick continued to work on the alignment throughout last week, making further progress. We can discuss this and any path forward to threading lasers into the TPC (which we should keep @Ross Corliss aware of) . We can also hear any report from @David Baranyai on the hallucination mitigation. Also, was the new firmware from @Finnelli, Kristina and @Kuczewski, John ever applied to the trigger boards ? Diffuse Laser: Once the gains have been declared balanced by @Thomas K Hemmick, we should resurrect the line laser and include it in the nightly data taking. As we do this, we should also take a TPC only run in global mode (BP w/ GL1) based on clock triggers with the diffuse laser firing and send that to @Kimelman, Benjamin for analysis. Its possible we may make another scan over intensity and work with @Azmoun, Bobto produce new response curves.
Contingency Box: @Mandracchia, Kevin has begun to look into the connector-izing of the boxes. We can discuss this with him. We should also make sure we take advantage of @shana Prifte 's offer to help. Analysis: @Luke Legnosky, @Go, Yeonju, and @osbornjd91 AT gmail.com all made analyses of the cosmic data with Luke and Yeonju producing gain measurements (see above). We can discuss any additional progress they want to report. Any other pressing business
Indico Link: https://indico.bnl.gov/event/27221/
-
[[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.17 at 9:30 AM ET,
Hughes, Charles [PHYSA], 03/16/2025
- Re: [[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.17 at 9:30 AM ET, Hughes, Charles [PHYSA], 03/16/2025
Archive powered by MHonArc 2.6.24.